Hi. Im going into 8th grade and my teacher gave us homework to do over the summer. I cant remember how to do any of these problems.
A little help is a ll i need,
Leg A : 12
Leg B: x-7
Leg C: x+1
WHAT DO I DO?

Answer by Theo(13342)

It appears this is a right triangle.
the formula for a right triangle is:
hypotenuse squared = leg 1 squared + leg 2 squared.
if leg A is the hypotenuse, then the formula would be:
(x-7)^2 + (x+1)^2 = 12^2
If leg B is the hypotenuse, then the formula would be:
12^2 + (x+1)^2 = (x-7)^2
if leg C is the hypotenuse, then the formula would be:
12^2 + (x-7)^2 = (x+1)^2